### v2

# Unsettling the Hegemony of Intention: Agonistic Image Generation

Abstract.
Current image generation paradigms prioritize actualizing user intention — “see what you intend” — but often neglect the sociopolitical dimensions of this process. As these tools become cornerstones of the media landscape, it is increasingly evident that image generation is political, contributing to broader social struggles over visual meaning. This sociopolitical aspect was highlighted by the March 2024 Gemini controversy, where Gemini faced criticism for inappropriately injecting demographic diversity into user prompts. Although the developers sought to redress image generation’s sociopolitical dimension by introducing diversity “corrections,” their opaque imposition of a standard for “diversity” ultimately proved counterproductive. In this paper, we present an alternative approach: an image generation interface designed to embrace open negotiation along the sociopolitical dimensions of image creation. Grounded in the principles of agonistic pluralism (from the Greek agon, meaning struggle), our interface actively engages users with competing visual interpretations of their prompts. Through a lab study with 29 participants, we evaluate our agonistic interface on its ability to facilitate reflection — engagement with other perspectives and challenging dominant assumptions — a core principle that underpins agonistic contestation. We compare it to three existing paradigms: a standard interface emulating current image generation tools, a Gemini-style interface that produces “diverse” images, and an intention-centric interface that suggests aestheticized prompt refinements before generation. Our findings demonstrate that the agonistic interface enhances reflection across multiple measures, but also that reflection depends on users perceiving the interface as both appropriate and empowering; introducing diversity without grounding it in relevant political contexts was perceived as inauthentic. Our results suggest that diversity and user intention should not be treated as opposing values to be balanced. Instead, interfaces can productively surface and navigate tensions between competing perspectives, enabling users to engage with and evolve their intentions in meaningful ways.
agonism, reflection, image generation interfaces
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Conference title.; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Interaction paradigmsccs: Human-centered computing Interaction design theory, concepts and paradigms
1. Introduction: Unsettling the Hegemony of Intention
What does “a Founding Father” look like? What does “Jesus” look like? What about “a migrant,” “a doctor,” “World War II”? The production and consumption of visual meanings is a sociopolitical endeavor. When people consider their mental image of “a Founding Father,” their preconceptions are in conversation with others’ interpretations about who a Founding Father is and — just as importantly — is not. Today, our collective “answers” to these questions increasingly draw from images that are generated by biased AI models and then proliferate in our digital media ecosystem (Jingnan, 2024; DiResta and Goldstein, 2024; Herrie et al., 2024). The political context of image generation was further underscored when in March 2024 Google’s Gemini model came under criticism for generating images with inaccurate racial and gender diversity — for example, depicting Nazi soldiers as people of color. Such examples show how the values embedded in image generation tools shape how we view and understand our social and political world. Despite the increasing politicization of image generation tools, the dominant paradigm behind most current AI image generation tools is the actualization of intention: they should enable anyone to see exactly what they were intending. This “intention is all you need” paradigm (Sarkar, 2024) neglects the political dimension of image generation. By inhibiting engagement with competing interpretations of user prompts, intention-centric paradigms run the risk of entrenching problematic user intentions (Cai et al., 2024).
Certainly, the Gemini controversy was as a meaningful attempt to grapple with the political dimensions of image generation by implementing corrections to address a lack of diversity in image generation models. Yet, while Gemini’s diversity intervention may have challenged the “‘hegemony of intention”, it did so in a way that caused significant public backlash, inciting “an explosion of negative commentary from figures such as Elon Musk who saw it as another front in the culture wars” (Milmo and Kern, 2024). The swift pushback indicates that top-down interventions are ill-suited to the dynamic political landscape that image generation tools find themselves situated in. In short, the problem is that shallow efforts that do not involve users in discourse but simply alter their intent can come across to users as patronizing.
In this work, we aim to build and evaluate an image generation interface that acknowledges the sensitive political contexts of image generation, including questions about diversity and representation, but unlike the Gemini case, encourages users to engage and reflect on their own intentions in relation to others on their own terms. Our design draws from agonistic pluralism (from the Greek agon meaning “struggle”), a political philosophy emphasizing the importance of conflict in the political realm and the productive ways it can be channeled instead of avoided. Given a user prompt, our image generation interface, Agonistic, collects and presents a range of interpretations of the prompt, prioritizing more controversial or challenging interpretations (Figure 1). Interpretations come with a justification grounded in concepts extracted from Wikipedia. Only after examining different interpretations and selecting one does a user then receive image generations.
A critical part of agonistic pluralism is reflection — the process of considering other viewpoints and challenging dominant assumptions. For this reason, we conduct a comparative lab evaluation of our interface for its ability to foster productive critical reflection. We build three other paradigmatic interfaces for comparison — Baseline: produces images generated from the user’s prompt with no further interaction; Diverse: rewrites the user’s prompt to be more diverse before image generation; and Reformulative: provides prompt reformulations to produce more aesthetic and preferable images, inspired by existing work in prompt reformulation. Twenty-nine participants used these interfaces to iteratively build and refine a collage of 10 images representing a human subject of their choice while thinking aloud.
Across several measures of reflection, we find that Agonistic induces the highest reflection out of any interface (§5), followed by Reformulative, then Diverse. For instance, when participants went to add images to a collage, they described the image as an expansion of their original intention 21% of the time while using Agonistic, but only 8%, 7%, and 3% of the time for Reformulative, Diverse, and Baseline, respectively. In addition, we find that participants express diversity as a justification for adding images to their collage $48\%$ of the time when using Agonistic, but only $28\%$ under Diverse. Our findings confirm that participants are likely to reject Gemini-style interventions due to a lack of factuality and authenticity, while Agonistic more authentically expresses diversity by situating user prompts and the generated images in relevant sociopolitical discourse. We find that getting users to acknowledge and prioritize diversity requires situating diversity in authentic political controversies, a different finding from current interventions that either seek to impose diversity for its own sake or constrain diversity with factuality. Agonistic aids critical reflection by giving users a sense of control and provide contextualizing elements to help them interpret images. These results point to richer ways that researchers and developers of AI tools can address ethical questions in generative AI. Instead of making top-down interventions, our work shows that AI tools can instead create space for open-ended contestation over values like user intention, factuality, and diversity. Ultimately, our findings highlight the importance of bringing HCI work to bear on AI ethics and provide more general design elements for building reflective AI systems.
2. Background and Context
2.1. Images as socio-political entities
Philosophers of meaning have argued that signs (things like words or images representing something other than themselves) can be “performed” to demonstrate membership within social categories (Butler, 1990; Hall, 1990, 1997) and invoked to effect changes in the world (Austin, 1956, 1962) — that is, signs do not just represent, but also play active roles in social life. Likewise, thinkers in history have challenged concepts of “objective” truth by calling attention to complex networks of competing narratives, power relations, and institutional practices (de Certeau, 1988; White, 1973) — what Michel Foucault calls “discourse” (Foucault, 1972) — that shape historical study. Thus, historical questions of “how things were really like” are far from straightforward because appealing to historical accuracy also requires considering ways that discourse shapes the historical record itself. Thinkers in media studies have further elaborated these insights in the specific context of visual media. Images are themselves political not only in the sense that interpretation of them is conditioned by cultural context (Barthes, 1981), but also in the sense that the act of creating or capturing an images always involves “imposing standards on [one’s] subject” (Sontag, 1977). For example, Laura Mulvey’s concept of the “male gaze” describes how the movement and position of the camera can frame women as passive objects of visual pleasure (Mulvey, 1975). These works suggest that all image production and consumption — including AI image generation — is implicated in social, political, and ethical issues: “individuals implicitly engage in ongoing struggles over visual dominance and its articulation with social formations through their involvement in image production, distribution, exposure, and consumption(Herrie et al., 2024). Our aim in this paper is to build an image generation interface which engages users in this complexity of image generation.
2.2. Agonistic Pluralism: A political framework for image generation
Given the political nature of image generation, what kind of political structures should be used to deal with this conflict? Some commentators have found the problem with the Gemini case to be that “tuning to ensure that Gemini showed a range of people failed to account for cases that should clearly not show a range(Milmo and Kern, 2024). If image generation is political, however, then there may not always be a clear division between cases that should or should not show diversity, but rather a pluralism of conflicting, sincerely held beliefs. For instance, though some reacted negatively to generated images of non-White Founding Fathers, public acclaim of the Broadway play Hamilton, which used people of color to portray Founding Fathers, implies that depicting historical figures in racially inaccurate ways is not necessarily always problematic. Consequently, we draw from democracy theory in this paper to inform our design because it offers a long tradition of scholarship aimed at answering questions of pluralism. Herrie et. al. suggest that within democracy theory, agonistic pluralism is apt for analyzing the political dimensions of image generation (Herrie et al., 2024). First developed by Chantal Mouffe, agonistic pluralism is a philosophy that sees struggles over meaning, like those in the Gemini case, as inherent and ineradicable parts of democratic society. The task for democracy then is not to eradicate such struggles, but to channel them into productive ends. Doing so requires at least two major conditions which we highlight here as relevant to our work. First, democracy should always allow for the possibility of political struggle by providing avenues for hegemonic views to be challenged. Second, Mouffe qualifies that agonistic pluralism also involves a basic level of respect for others: in contrast to purely antagonistic struggle in which opponents view each other as enemies to be destroyed, agonistic struggle requires viewing opponents legitimate adversaries “whose ideas we combat but whose right to defend those ideas we do not put into question(Mouffe, 2000).
2.3. Current values in image generation interfaces
Many current works aim to help users produce “better” images without necessitating additional user interaction, invoking the value of intention actualization. Prompt reformulation methods automatically rewrite prompts to produce more aesthetic or generally preferred (Zhan et al., 2024; Xu et al., 2024; Datta et al., 2024; Chen et al., 2024) images, motivated by the goal of “convey[ing] the user’s intended ideas(Mo et al., 2024) or “preserving the original user intentions” (Hao et al., 2024) Other work builds interfaces to help users explore and iterate over prompts, invoking similar values: “‘align the user’s creative intentions with the model’s generation”  (Wang et al., 2024b); “align with [users’] intended creative output” (Brade et al., 2023). This emphasis on actualizing user intentions is described by Sarkar (2024) as the “intention is all you need” paradigm: “Contrary to the assumption that GenAI merely executes human intentions, it also shapes them […] induc[ing] ‘mechanised convergence,’ homogenising creative output, and reducing diversity in thought.” Cai et al. (2024) characterizes interfaces that overly emphasize cognitive ease of use as “easy to learn at the expense of the power and complexity necessary to do hard but valuable work or learn uncomfortable truths”. In response to these critiques, our work explores how users might question and revise their intentions, for example by reflecting upon limiting assumptions.
Another notable value in image generation is diversity. Image generation models have well-documented issues with diversity, often failing to appropriately represent demographic minorities and resorting to stereotypes (Ali et al., 2024; Wang et al., 2023, 2024a; Chauhan et al., 2024; Naik and Nushi, 2023, inter alia). Approaches to address diversity issues include finetuning the model with some diversity objective (Miao et al., 2024; Esposito et al., 2023), intervening on internal model states (Friedrich et al., 2023; Zhang et al., 2023; Zameshina et al., 2023; Chuang et al., 2023), or rewriting the prompt to emphasize diversity (Microsoft, 2024). Our work takes another angle on diversity, aiming to engage users with competing visual interpretations of a subject rather than directly generating “diverse” images. We further discuss this issue in §7.1.
2.4. Agonistic design and reflection in HCI
In contrast to this intention-centric paradigm, Kate Crawford poses the question: “What might happen if we brought a model of agonism to understanding algorithms?(Crawford, 2016). She answers that agonistic design requires acknowledging that “complex, shifting negotiations are occurring between people, algorithms, and institutions, always acting in relation to each other(Crawford, 2016). These suggestions are in line with more recent literature in support of design principles that similarly seek to disrupt rather than reproduce user intention (Cai et al., 2024; Sarkar, 2024; Herrie et al., 2024). As applied to image generation, agonistic design entails that developers might avoid imposing a single standard on image generation outputs in favor of bringing controversy to users’ attention and providing “resources for more people to participate in that dissensus(Crawford, 2016).
Agonistic design also finds affinities in existing HCI work on designing for reflection, understood here as “intellectual and affective activities in which individuals engage to explore their experiences in order to lead to new understandings and appreciations(David Boud, 1985). Based on a review of HCI research about reflection, Fleck and Fitzpatrick set forth a five-level framework of reflection, in which each level is distinguished from the last by increasingly deep consideration of different views. The last stage of reflection, called critical reflection, involves challenging personal assumptions by considering a wide range of social and ethical issues (Fleck and Fitzpatrick, 2010). Halbert and Nathan further elaborate this last stage of critical reflection as the result of a two-stage process in which users first encounter a discomforting experience, and then become critically aware of their own biases by working through discomforting feelings, rather than avoiding them (Halbert and Nathan, 2015). Though comparatively little HCI research on reflection has been done for art and culture domains (Bentvelzen et al., 2022), we aim to fill that gap by designing image generation interfaces for user reflection in this paper.
3. Four Paradigmatic Interfaces for Image Generation Interactions
[…]
(a)  Baseline
[…]
(c)  Reformulative
[…]
(b)  Diverse
[…]
(d)  Agonistic
Figure 1. Annotated screenshots from each of the paradigmatic image generation interfaces evaluated in our study. See §3 for details.
To holistically evaluate how agonistic design might encourages reflection (§2.4), we create four image generation interfaces: Baseline, which accepts the user’s prompt and outputs generated images with no further interaction, and three paradigmatic interfaces each embodying a dominant approach or value — Diverse for diversity, Reformulative for intention actualization, and Agonistic for agonism. In this section, we describe the important frontend and backend elements of each interface, with more details in Appendix §B. Annotated screenshots are displayed in Figure 1.
Baseline  $\cdot$  We design a baseline interface as a control for the study. In this interface, users enter a prompt and are displayed four images generated from the prompt. We use a lightweight open-source image generation model (Black Forest Labs, flux-schnell) throughout all interfaces in the study, to avoid relying on interfaces that perform under-the-hood prompt rewriting (like the DALL-E API).
Diverse  $\cdot$  Diverse represents the paradigm of image generation that explicitly corrects for diversity issues by rewriting prompts in more diverse ways, as was reported in the Gemini case. To create this interface, we use a leaked alleged DALL-E prompt as reported by (Milmo and Kern, 2024), which instructs the model to “diversify depictions of ALL images with people to include DESCENT and GENDER for EACH person using direct term.” We instruct GPT-3.5 to rewrite the user prompt with the above instructions and use the rewritten prompt to generate four images that are then displayed to the user. Although we recognize that the prompt may not match the exact approach used in the Gemini case, it anecdotally yields similar results to reported images for a variety of historical prompts. Moreover, this approach makes Diverse a realistic interface to compare against, since it uses the (alleged) prompt for the DALL-E API, which is widely in use.
Reformulative  $\cdot$  The Reformulative interface embodies the value of intention actualization and is inspired by §2.3; it supports users by producing prompt reformulations that add detail to the prompt (e.g., adding useful semantics or stylistic indicators) in ways that are likely to result in more aesthetic or preferred images. This interface is similar to interfaces like Promptify, which allows users to iteratively refine their prompts based on a variety of AI-generated reformulations (Brade et al., 2023). Skilton and Cardinal propose using AI to generate a list of “suggested descriptors” that the user can then choose from and modify to create images that depart from stereotypical representations (Skilton and Cardinal, 2024). We instruct GPT-4o to generate a diverse set of detail-adding reformulations of the user prompt with in-context examples from or inspired by Brade et al. To aid visual navigation of reformulations, we generate a sample image (“thumbnail”) for each reformulation and display both in a list of suggested reformulations to the user. The user can choose a suggestion and modify the suggestion before using it to generate images. This interface is designed to have a comparable interaction level and similar features where applicable to Agonistic for fair comparison.
Agonistic  $\cdot$  The Agonistic interface attempts to forefront ambiguities and controversies over visual interpretation to the user, drawing from §2.2 and §2.4. We implement Agonistic by creating a multi-step retrieval-augmented workflow that researches controversies about the user prompt from Wikipedia and presents them to the user. Our work builds on insights from Cox et. al., who find that presenting users with maximally distant phrases (measured by language embeddings) is effective at enhancing diversity in creative ideation (Cox et al., 2021). Because Wikipedia allows public contributions, we believe it is more likely to capture relevant prompt controversies. $\cdot$  We identify the main subject of the prompt with GPT-3.5 and perform a Wikipedia search for 50 pages related to the main subject. We then use GPT-4o to filter for 40 pages relevant to the user prompt based on their titles (excluding, for example, soccer player Gabriel Jesus from a search for “Jesus”). Next, we compute a controversy score for each page and rank pages by controversy, selecting the top 20 most controversial pages. The controversy score is calculated by dividing the total number of edits by the number of unique editors, following findings from Kittur et. al. that controversy is positively correlated with total number of edits and negatively correlated with the number of unique editors (Kittur et al., 2007). We chose this particular controversy metric for its efficiency and strong qualitative performance.  $\cdot$  We then instruct GPT-4o to generate 5 mental images an “average person” might have of the main subject and provide this in-context to the interpretation generation call. For each page, GPT-4o then selects 4 sections based on their titles to read and produces an interpretation with 4 fields: 1) a section summary explaining the main points of the cited page section; 2) a description of what the user’s main subject looks like, taking into account the user’s full prompt; 3) a source with the referenced page and section; 4) a justification explaining how the section content justifies the description. The section summary is not shown to the user but is generated first to reduce model hallucinations.  $\cdot$  We instruct GPT-4o to generate descriptions that challenge the provided mental images, and phrase justifications in the form “you may assume…, but… .” These decisions were made to increase the likelihood that users would encounter discomforting suggestions, creating opportunities for critical reflection. Finally, like in Reformulative, we generate a “thumbnail” for each interpretation. Users are shown a list of interpretations with descriptions, thumbnails, and sources; when users click on an interpretation, the card expands and the justification becomes visible. After a 3-second wait period to encourage users to read the justification, the user can “accept” the interpretation and generate images, as well as edit the description text and re-generate. The design iteration of this interface is documented in §B.1.
4. Experimental Design
To evaluate all four interfaces with respect to reflection, we conduct a within-subjects comparative lab study and ask (RQ1): How each interface induces reflection using a variety of measures for user reflection; and (RQ2): Why each interface induces reflection by examining how other variables could explain the reflection observations in RQ1.
4.1. User Task and Interview Structure
Each lab study session was conducted 1-on-1, lasted around one hour, and consisted of the setup and user task.
Setup ($\sim$15 mins). Each participant was randomly assigned to an ordering of interfaces, with Baseline fixed as the first interface (e.g., [Baseline Agonistic, Diverse, Reformulative] — 6 unique orderings), to account for learning effects between non-Baseline interfaces. The interviewer received consent to record the interview and use anonymized data. Then, to familiarize the participants with the interfaces, participants were guided through using each interface in their assigned order with the prompt “a person.” Afterwards, participants selected a subject (prompt) for their task; they were told to choose one of interest to them and encouraged to choose one in a randomly pre-assigned category. Subjects all featured people and were divided into three categories — identity / demographics, history, and political issues. This choice of categories allowed for potential reflection on issues of visual representation. Participants were provided with examples in each category for inspiration but also allowed come up with their own subjects. See §A.1 for participants’ chosen prompts and §A.2 for example prompts.
User Task ($\sim$45 mins). After setup, participants were informed of their task: to create a collage of ten images which represented all relevant aspects of the subject. This task forced participants to make nuanced choices about what to include and exclude in visual representation of their chosen subject. Participants constructed an initial collage of ten images with Baseline. Then, they improved their collage by interacting with the other interfaces (Diverse, Reformulative, Agonistic) in their assigned order. If participants produced an image they wanted to add to their collage, they picked one of their existing images to replace. The task design of accumulating collage improvement gave permanence to users’ decisions about visual representation, provided a holistic picture of how users interacted with interfaces across different stages of exploration, and was a more natural and engaging task for participants compared with other experiment designs tested. Participants were instructed to “think out loud,” commenting on their choices to include or exclude images in their collage. Figure 8 displays an example collage progression from the study.
4.2. Recruiting Procedure and Study Chronology
After receiving IRB approval from our university, we sent recruiting materials to undergraduate and graduate students in university departments and student groups via online communication channels. We aimed to represent a variety of backgrounds, experiences, and interests in our participant pool. We piloted an initial version of our study with 11 individuals and revised the task to its current form (as described in §4.1) in response to time constraints and perceived confusion about the task. Then, the two co-first authors individually conducted interviews with 29 individuals over Zoom. Participants were compensated \$$\$20$ per hour. More information about participants is available in §A.1.
4.3. Measures
Given the work discussed in §2.4, we measured reflection by looking quantitative and qualitative signals of how much individuals questioned or changed prior assumptions after using our interfaces. While it is more common to use purely qualitative metrics when measuring reflection (Bentvelzen et al., 2022), we believed a mixed-methods approach offered unique benefits for our study because quantitative metrics would allow us to more precisely compare reflection across interfaces. We drew from common metrics in mixed methods studies of reflection such as interviews, questionnaires, and self-reports (Bentvelzen et al., 2022).
Survey. After participants interacted with each interface, they provided responses on a 5-point Likert-style scale to measure perceived rethinking (how much their mental image changed), satisfaction, appropriateness, and control (see §A.3). For the Reformulative and Agonistic interfaces, we also collected responses about how interesting the suggestions/interpretations were. Quantitatively measuring self-reported rethinking helped capture reflection not have been recorded by other artifacts like collages (for instance, when an image caused reflection but was not added). We measured non-reflection variables (e.g., control) to explain why interfaces might have different reflection.
Interview Coding. After conducing interviews, the two co-first authors coded interview transcripts to systematically extract further information about user experience. They independently coded a small subset of interviews before coming together to build a coding ontology. We divided codes into intents (how user intents changed while adding an image, roughly correlated to different degrees of reflection) and values (reasons users give for adding an image). Values are not mutually exclusive, because a user can invoke multiple values when adding an image. The two first co-authors then independently coded a subset of three interviews to compute inter-rater reliability (IRR), revising the coding ontology and independently re-coding interviews as needed to resolve disagreements. The final IRR was 0.67 based on a weighted average of Cohen’s Kappa scores across value codes. After coding all interviews, the two co-first authors reviewed all images marked with non-Direct intent together to establish more agreement on intent codes. See §C for our IRR methodology and §5.2, §6.2 for a delination of intents and values.
5. How do interfaces induce reflection?
(a) Overall, by interface.
(b) Breakdown within interface by prompt categories: “H” for History, “P” for Politics, “I” for identity.
Figure 2. Participant ratings on a 5-point scale for how much interacting with the interface made them rethink their mental picture.
In this section, we explore how different interfaces induce reflection by investigating variables that directly measure some aspect of reflection collected in the study. The aim is to paint a holistic picture of how reflection occurred under different interfaces. In Table 1 we provide two examples of reflection for each of the interfaces from our interviews. To characterize reflection across different interfaces, we investigate users’ self-reported change in mental image (§5.1) and reflection on a per-image level (§5.2) as described in §4.3.
PID Examples of Participant Reflection
Baseline P24 The participant noticed that all the images for “the constitutional convention” featured the delegates sitting. In subsequent generations, they specified that the delegates should be standing. They reflected: “when I kept generating those prompts and they were sitting […] I don’t think I would have specified whether they were sitting or standing before. Since I thought about it, I had to… play around with the prompt to get them to stand”. They further remarked: “Interacting with this interface […] made me pay more detailed attention to what I perceive would have actually happened at that historical event.
P4 After generating a few images for the subject “a queer community” which seemed to “rely on heavy stereotypes […] like pride parades”, the participant wanted to generate “everyday examples of queer people”. After producing a few images, they remarked “now you lose the queer part.” After a few moments, they reflected: “Well now you really got me questioning myself […] what is a queer community? Do they all have to be wearing pride flags? How do you tell? Now I’m getting in my own head about it.
Diverse  P19 The participant was attempting to generate an image of “A Syrian refugee working in Germany” when Diverse generated an image of a woman in a healthcare setting. The participant paused and remarked, “can you train to be a doctor [in] 12 years?” They later explained that they were unsure whether a Syrian refugee would have been able to study long enough to become a licensed medical professional since leaving Syria, but reasoned that “the situation has gone on long enough [that] they have studied and integrated in the societies they’re in, where [they’re] able to actually finish something in the medical sciences.
P29 The participant had previously generated all images of White people for the prompt “a Jewish person” using Baseline. After Diverse produced images of Black people, the participant felt that the images were not appropriate but became confused upon further reflection: “Okay, I’m focused on the Black […] person here because, this is because my lack of knowledge […] I don’t know what makes a person Jewish other than the religious element [but…] to be Jewish, I think you have to be born in a family […] I could be wrong.
Reformulative P10 With the prompt “A gun owner,” Reformulative generated the suggestion, “An elderly man polishing his grandfather’s antique shotgun.” Upon reading the suggestion, the participant reacted, “I didn’t even think about that kind [of] collector gun owner,” before using the suggestion to generate and add an image to their collage. They later reflected that the interface “reminded [them] of things [they] forgot about” and “sparked a new mental image” for them.
P19 Given the prompt “A gun control advocate,” Reformulative generated a suggestion of “A gun control advocate engaging in a community discussion.” The participant reacted, “I like that it’s suggested positioning the advocate in a community space, where all of the other images [from previous interfaces] were very portrait-style.” They used the suggestion to generate and add an image to their collage, later sharing that they liked how Reformulative helped them “think of things that [they] wouldn’t think to ask.
Agonistic P5 After interacting with Agonistic on the subject “World War II” which showed many less well-known and more local actualizations of WWII, the participant (whose family was strongly impacted by WWII) revised their design statement from “WWII involved mass conflicts that…” to “WWII involved conflicts that…”. They reflected: “I think it just involves conflicts in general now. It’s making me rethink that word [mass]: history tends to focus on real large-scale things, but I find that all of them are important to me personally. I feel like it’s actually made me reconnect to the way I think about World War II. And instead of trying to project what I would like, think that other people need to see about World War II.
P28 The participant initially described their mental image of the prompt “the signing of the Declaration of Independence” as “A group of […] white men like Thomas Jefferson standing around a table signing a document.” When Agonistic generated images of non-White people, drawing from sources like “Haitian Declaration of Independence,” the participant observed, ”when […] I put Declaration of Independence, I assume a mental image of the US Declaration of Independence, but so many nations have Declarations of Independence.” They continued, “It makes me realize […] this assumption of bias… Oh my gosh, I was so narrow-minded in my approach.
Table 1. Two examples of reflection observed during our interviews when using each interface.
5.1. Change in mental image
The change in mental image is reported by participants after they use each interface and is measured by agreement with the statement “Interacting with the interface made me rethink the mental picture of the subject I had right before using this interface” on a 5-point Likert-style scale. We take higher reported changes in mental image to be indicative of more reflection. Averages are reported in Figure 2 and Table 4.
Agonistic induces the most reflection as measured by self-reported rethinking. Participants’ mental images change more under Agonistic than Reformulative ($p\ll 0.01$), and more under Reformulative than Diverse ($p=0.05$). Neither Reformulative nor Diverse surpass the Baseline(likely due to the fact that participants interact with Baseline first), but Agonistic does ($p\ll 0.01$). The ‘hierarchy’ of interfaces by reflection is thus $\left[\textsf{{Agonistic}}>\{\textsf{{Baseline}},\textsf{{Reformulative}}\}>\textsf{{Diverse}}\right]$. A similar ‘hierarchy’ emerges when considering how participants rank interfaces by how much their mental image changed retroactively at the end of the study (see §D.1), as well as measuring how much participants’ “design statements” changed after using each interface (see §D.3). Agonistic further induces greater unique rethinking based on an analysis of ordering effects (see §D.2).
Rethinking under Agonistic is robust to prompt category. Though Diverse and Reformulative see significant drops in rethinking for historical prompts compared to political prompts ($-0.27$ and $-1.27$, respectively) the same drop is not observed for Agonistic ($-0.00$). While all domains have roughly similar rethinking under Agonistic, political subjects lend themselves to more rethinking overall under Diverse and Reformulative than historical subjects.
5.2. Reflection when adding images
Next, we aim to capture a richer picture of reflection by analyzing the distribution of intent codes across interfaces. As explained in §4.3, intents code for how user intents changed while adding an image to their collage. Intent codes are defined as follows: 1) Direct: the user already intended to generate the image; 2) Reminder: the user is reminded of a detail about the prompt they already accepted but forgot; 3) Expansion: the user accepts an image as a different valid interpretation of the prompt than their original intent; 4) Challenge: the user realizes an error or otherwise experiences a significant change in their original intent. Direct intent represents little to no reflection, whereas Reminder, Expansion, and Challenge represent increasing degrees of reflection. The distribution of intents across interfaces is shown in Fig. 4
Agonistic induces the most reflection as measured by quantitative interview analysis. Participants replace the most images on average with Agonistic(3.1), followed by Reformulative (2.9) and Diverse (1.7). We observe Direct intent most frequently with Baseline ($98\%$), followed by Diverse ($95\%$), Reformulative ($90\%$), and a sharp drop in Agonistic ($67\%$). Agonistic also has the highest proportion of Reminder and Expansion of intent across all interfaces ($11\%$ and $19\%$ respectively, compared to the next-highest $4\%$ and $6\%$ for Reformulative), and is the only interface to Challenge  intent ($4\%$). An example of Challenge intent by P28 is given in Table 1. A minor difference with §5.1 is the finding that Diverse induces more frequent reflection than Baseline (see §D.6).
Figure 3. Breakdown of how user intents changed when adding images for each interface (discussed in‘§5.2; Table 7).
Figure 4. Mean responses on a 5-point scale for Rethink, Appropriateness, Control, and Interest (discussed in §6.1; Table 5).
6. Why do different interfaces induce reflection?
To explain the observed differences in reflection under different interfaces as discussed in §5, we look at the relationship between reflection and other variables: perceived interface-level properties like appropriateness and control (§6.1), participants’ values when adding images (§6.2) and why they reject images (§6.3), and qualitative themes (§6.4).
6.1. Perceived Interface-Level Properties
Reformulative and Agonistic improve upon Baseline and Diverse in appropriateness and control ($p\leq 0.05$ for all). Indeed, in general, there is a weak but significant correlation between rethinking and both appropriateness ($r=0.22$, $p=0.02$) and control ($r=0.27$, $p=0.003$) in general. Agonistic is ranked above Diverse by $81\%$ of participants for appropriateness and $86\%$ for control; Reformulative by $85\%$ for appropriateness and $92\%$ for control. Although Diverse cannot be distinguished significantly from Baseline by appropriateness, it can by control — Diverse has lower control than Baseline ($p\ll 0.01$), meaning that rewriting user prompts to encourage diversity results in a noticeable feeling of decreased control among participants. That being said, Agonistic and Reformulative are not themselves distinguishable by control or appropriateness ratings ($p\gg 0.05$). Additionally, participants find Agonistic’s possible interpretations more interesting than Reformulative’s suggested reformulations ($p\ll 0.05$).
6.2. Image-Level Values: Why users add images
Figure 5. Distribution of values cited for images added with each interface (discussed in §6.2; Table 6).
Figure 6. Breakdown of intents by value cited when adding image (discussed in §6.2); Table 8.
We also explore whether different values may serve as a relevant explanatory variable for types of reflection (intents). Values code for why participants add images, or in other words the reasons that participants invoke when selecting images. The values in our coding ontology are as follows: 1) Realism: the image represents how the user thinks the world actually is; 2) Familiarity: the image fits the user’s assumptions, regardless of whether the user thinks they are representative of how the world actually is; 3) Diversity: the image portrays an underrepresented aspect of the prompt that the user believes is normatively important to include; 4) Aesthetics: the user otherwise likes how the image looks. The distribution of values across interfaces and intents is shown in Fig. 6 and 6.
Agonistic is the only interface for which Diversity is invoked more frequently than Familiarity. In all interfaces except Agonistic, Familiarity is invoked most frequently, followed by Realism, Diversity, and Aesthetics. In contrast, for Agonistic, participants invoke Diversity most frequently, followed by Realism, Familiarity, and Aesthetics. Interestingly, compared to Baseline, participants invoke Diversity less frequently for Diverse ($-2\%$), but more frequently for Reformulative ($+12\%$) and Agonistic ($+19\%$).
Diversity is associated with the highest rate of reflection. The majority of images across values are added with Direct intent. Among values, Aesthetics  is associated with the highest highest proportion of Direct intent ($96\%$), followed by Familiarity($95\%$), Realism ($89\%$), and Diversity ($75\%$). Diversity is associated with the highest degrees of reflection, with $7\%$ of images added with Reminder, $16\%$ of images added with Expansion, and $2\%$ of images added with Challenge. Since Agonistic also has the highest proportion of Diversity values, these findings suggest that Diversity serves as a strong explanatory variable for reflection.
6.3. Image-Level Values: Why users reject images
We carry out a qualitative analysis of image rejection in lieu of a quantitative analysis because the unit of rejection is not as well-defined as adding images.
Participants reject images due to factual inaccuracies, especially when interacting with Diverse. P7 said that racially diverse pictures of ‘The Chinese Communist Revolution’ “looked like modern people were slapped into the wrong time period,” and P28 experienced “pushback” against racially diverse pictures of “The Signing of the Declaration of the Independence.” In other interfaces, participants similarly cited lack of factuality as a reason for rejecting images. For instance, P2 reacted “Middle class people don’t dine out very often, and especially not at fancy places,” when Baseline generated an image of people at an upscale restaurant for the prompt “A middle class person.”
Participants also reject images because of an excessive adherence to factuality. P26, generating images of “A Syrian refugee” noted that Agonistic “presented more of what I know to be true, […] the harsh reality of it. […] The other interfaces […] were more instrumental in helping me expand beyond that.P7 shared that they preferred Reformulative over Agonistic because Reformulative did a better job “highlighting all different aspects of [their prompt], instead of just the realism images.” Such comments suggest that users do value an some level of diversity in generated images that may extend beyond what they consider to be strictly factual.
Participants reject images due to lack of familiarity, even where they acknowledge that an image may be factual. When using Agonistic, participants may believe interpretations to be factual (being extracted from Wikipedia) but still reject them because they are unfamiliar. For example, with the prompt “An Arab person”, P25 reacted to an image of a Sudanese Arab person from Agonistic by remarking, “Oh, Sudanese Arab […] I think that’s an identity that I haven’t represented here that I would like to.” They later decided to reject the image by providing the following reasoning: “It just does not resonate with me. I guess I’m also not familiar with Sudanese Arabs as much.
6.4. Qualitative Themes
We present several themes from qualitative analysis further explaining the differences in reflection observed in §5.
Text helps users contextualize, interpret, and decide on images. Throughout the user studies, we find that participants prefer interfaces that help them contextualize and interpret images (Reformulative and Agonistic over Baseline and Diverse), even though ultimately they are only adding the images (and not the accompanying text) to the collage. Firstly, contextualizing content is cognitively stimulating: “[the] suggestion feature… by nature makes your brain think about alternatives” (P24); it “I wouldn’t have thought of [these perspectives] without the text” (P1). Participants feel that it is meaningful and important to interpret the image, which is mediated by thinking about the relevant text. P1 said that the generated images “adhere to the prompt but you have to figure out the image yourself”. P15 said of Reformulative: “I like with this interface. It gives me a lot more of an explanation. It kind of tells a story that I might not know just by looking at it… I think that it could help me kind of change what I’m looking for.
‘Authenticity’ situates diversity in an engaged political context. Although both Diverse and Agonistic end up displaying demographically diverse images to users, we find overwhelmingly that participants prefer the type of diversity offered by Agonistic to that offered by Diverse due to a greater perception of authenticity. Diverseis not only perceived as factually inaccurate (§6.3), but nontransparent, alienating, and ungenuine. P17 said they couldn’t “get a clear idea from anything” from the interface. P5 reacted when racially and gender diverse World War II soldiers generated by Diverse: “I find them not just […] historically inaccurate, I think it gives narratives that some people will not dig deeper into, and […] can lead to a lot of misconceptions down the line of colorblind racism.” In contrast to Diverse, participants report that the perceived authenticity of Agonistic engages them in the political context of image generation (§2.2). P10 liked that Agonistic was “more grounded in reality” — “you’re like, ‘yeah, real people exist.’P28 described Agonistic: “[it] changed my perception of what the prompt itself was […] but [Diverse] feels like it’s […] trying to change our perception of the event itself that I was already thinking of […Agonistic] tries to bring in diverse perspectives and increase representation of other groups […] in a way that felt more authentic and can change people’s assumptions in a healthier way.
7. Discussion
7.1. The inefficacy of inauthentic diversity
Though it is common to conceive of diversity and factuality as conflicting values, our findings suggest that context is more important for authentic diversity. Some participants preferred Diverse to Baseline because they felt that the demographic diversity of Diverse was appropriate and either more true to representing the subject or more familiar to their experience. For example, in response to images of a diverse crowd worshiping Jesus, P6 said,“It’s nice having […] diverse people, [be]cause I think that’s just how it is.” Nonetheless, as discussed in §6.3, many participants echoed concerns about conflicts between diversity and factuality — diversity’s “factuality tax” (Wang et al., 2023) — that also surfaced in the online reaction to the Gemini case (Milmo and Kern, 2024; Robertson, 2024). Yet notions of factuality are seldom so straightforward (§2.1). For example, when creating a collage for “the women’s suffrage movement”, P18 was hesitant to include images of non-White women because many American suffragette leaders espoused racist views, but changed their mind after engaging with an interpretation from Agonistic highlighting the role that Black women played in leading their own suffragette organizations. Here, adding the image did not obscure the facts of historical racism but highlighted other relevant aspects in the complex history of the suffragette movement. Thus although the “factuality tax” is perceived as a necessary cost of diversity, it is not clear that image generation must always be thought of along a strict dichotomy between factuality and diversity. Given that images are situated in a social context, developers should turn attention from a narrow focus on generating only the right kinds of images to also generating the right kinds of conte(x/n)t — something that users in our study found valuable (§6.4) — which can give rise to new ways of seeing and understanding the prompt.
Several participants bring up aspects of a less discussed angle with Diverse’s approach to diversity: its authenticity. Taking a definition that “to say that something is authentic is to say that it is what it professes to be, or what it is reputed to be, in origin or authorship” (Varga and Guignon, 2023), Diverse can be critiqued not only as unfactual but also insufficiently aligned with the underlying goals of diversity—such as respecting and representing all groups of people. In §6.4, we documented how situating diversity in a political context encourages user perception of authenticity, as further exemplified by P18’s reflection about the suffragette movement mentioned above. These findings support the idea that the “factuality tax” is an overly restrictive framework to conceptualize authentic diversity, which should be understood neither in terms of diversity for the sake of diversity nor diversity rigidly constrained by factuality. Moreover, our findings that participants invoke Diversity at higher rates when adding images with Agonistic  ($45\%$ of images) than Diverse ($24\%$ of images) suggest that agonistic pluralism offers a superior framework for actualizing authentic diversity by engaging the user with relevant political context.
7.2. Engaging with users experientially
Our study suggests that something is missing in a picture in which user engagement with interfaces is driven by the interface’s factuality (realism, consistency) and diversity: the role of familiarity. As discussed in §6.2 and §6.3, many images are accepted and rejected because they are un/familiar rather than because they are un/factual or un/diverse. If we want to steer users towards desired ideals like critical reflection, interfaces should treat them not solely as judges for factuality and diversity, but as people who engage with images on the basis of familiarity and experience. The design of Reformulativeleans into giving users aestheticized and familiar experiences, possibly at the expense of reflection: “ if there was a change that I wanted to implement in these pictures, but I just didn’t have the words for it, I think [Reformulative]… gave the options. I didn’t have to… think about what I was looking for. I could just see examples of what I was looking for” (P19). Meanwhile, Agonistic focused not only on producing factual and diverse outputs via Wikipedia-grounded interpretation sources, but also in explicitly engaging with and challenging users’ experience and sense of familiarity. Our findings therefore further illuminate how core principles of agonistic pluralism, which emphasizes the “crucial role played by passions and affects in securing allegiance to democratic values.” (Mouffe, 2000).
7.3. Limitations and Future Work
Interface-level. Wikipedia entries are overtly sites of discourse and negotiation, and prone to political, cultural, and other forms of bias (Hube, 2017; Umarova and Mustafaraj, 2019). Even though our discovery process attempts to find varying views on subjects, this bias may filter into the process in a way that hinders agonistic user engagement. Future work might better understand the way that knowledge communities (e.g. Wikipedia, Britannica (Greenstein and Zhu, 2014)) connect with user interfaces that seek to expose users to the social and discursive processes (e.g., Wikipedia page edit history) by which information is produced. Future work on agonistic principles in generative AI interfaces might consider more explicit and iterative representations of discourse (e.g., graph-based) rather than Agonistic’s summarizing and linear presentation of discourse/interpretations.
Study-level. Although we tried to mitigate these factors, our study may be prone to novelty and participant response bias (Dell et al., 2012). Our study favored lower scale (29 participants) abut more high-resolution (1-hour interviews, collecting self-reported measures, task artifacts, qualitative responses, etc.) examinations of user experiences. Additionally, our sample population was recruited from university students in the United States, all from a single university. Future work might systematically examine how larger-scale populations engage with agonism in generative (image) AI interfaces.
8. Coda: Bricks and Windows
We introduced agonistic pluralism as a challenge to the ‘hegemony of intention’ in image generation interface design, demonstrating that an interface built on agonistic principles engaged individuals with competing interpretations in the salient discourse, making them more critically reflective. Gilles Deleuze and Felix Guattari delightfully noted that “A concept is a brick. It can be used to build a courthouse of reason. Or it can be thrown through the window” (Deleuze and Guattari, 1987). Instead of obsessing over “structurally sound” interfaces that perfectly embody capital-I Intention, capital-F Factuality, and/or capital-D Diversity, maybe our interfaces should give users some bricks to throw at the windows — they were always brittle anyways, bordering on stale, ready to be broken. Perhaps the view outside will be clearer.
Appendix A Full Study Materials and Information
A.1. Participant Backgrounds
Participants came from the following departments: Computer Science, Physics, Math, Economics, Informatics, Comparative History of Ideas, English, History, Business, American Ethnic Studies, Law, Journalism, Education, Library and Information Sciences, Gender Women and Sexuality Studies, Communications, Cinema and Media Studies, and Political Science. Out of the 29 participants in the study, 21 were undergraduate students and 8 were graduate students.
PID Selected Prompt Category Interface Order
P1 A Korean person Identity Diverse, Reformulative, Agonistic
P2 A middle-class person Identity Agonistic, Diverse, Reformulative
P3 A designer Identity Reformulative, Diverse, Agonistic
P4 A community of queer people Identity Agonistic, Reformulative, Diverse
P5 World War II Politics Reformulative, Agonistic, Diverse
P6 Jesus History Diverse, Agonistic, Reformulative
P7 The Chinese Communist Revolution History Diverse, Agonistic, Reformulative
P8 An Asian person Identity Agonistic, Diverse, Reformulative
P9 An immigrant Identity Agonistic, Reformulative, Diverse
P10 A gun owner Politics Diverse, Reformulative, Agonistic
P11 The French Revolution History Agonistic, Reformulative, Diverse
P12 A fascist Politics Reformulative, Diverse, Agonistic
P13 A socialist Politics Diverse, Agonistic, Reformulative
P14 The Israel-Palestine Conflict Politics Reformulative, Agonistic, Diverse
P15 A mass shooter Politics Diverse, Reformulative, Agonistic
P16 A gun control advocate Politics Reformulative, Diverse, Agonistic
P17 World War II History Agonistic, Diverse, Reformulative
P18 The women’s suffrage movement History Agonistic, Diverse, Reformulative
P19 A Kamala supporter Politics Reformulative, Diverse, Agonistic
P20 Lord Krishna History Diverse, Reformulative, Agonistic
P21 An activist Identity Agonistic, Reformulative, Diverse
P22 A Democrat Politics Reformulative, Agonistic, Diverse
P23 A Roman gladiator History Agonistic, Diverse, Reformulative
P24 The Constitutional Convention History Reformulative, Agonistic, Diverse
P25 An Arab person Identity Reformulative, Diverse, Agonistic
P26 A Syrian refugee Politics Diverse, Reformulative, Agonistic
P27 A Christian person Identity Diverse, Agonistic, Reformulative
P28 Declaration of Independence Signing History Agonistic, Reformulative, Diverse
P29 A Jewish person Identity Diverse, Agonistic, Reformulative
Table 2. Participant IDs with their selected prompt, category, and the order in which they encountered each of the non-Baseline interfaces.
A.2. Full Topic List
Participants were provided the following list of example topics/subjects within each of the three categories.
  • $\bullet$
    Identity & Demographics
    • $\circ$
      Race: “a Black person”, “a White person”, “an Asian person”
    • $\circ$
      Ethnicity: “a Hispanic person”, “a Jewish person”, “an Arab person”
    • $\circ$
      Nationality: “an American person”, “a Chinese person”, “a Brazilian person”, “a Swedish person”, “a South African person”
    • $\circ$
      Class: “a working class person”, “a middle class person”, “a member of the elite”
    • $\circ$
      Family history: “an immigrant”, “a refugee”, “an indigenous person”
    • $\circ$
      Occupation: “a nurse”, “a student”, “a businessperson”, “a CEO”, “a/the Pope”
    • $\circ$
      Religion: “a Christian person”, “a Muslim person”, “a Buddhist person”
    • $\circ$
      LGBTQ+: “a transgender person”, “two men getting married”
  • $\bullet$
    History
    • $\circ$
      Historical people: “a Founding Father”, “a Nazi soldier”, “Jesus”, “the prophet Muhammad”, “a Roman gladiator”
    • $\circ$
      Historical events: “World War I”, “the French Revolution”, “the Holocaust”, “the Great Depression”, “the American Civil War”, “the Chinese Communist Revolution”, “the fall of the Berlin Wall”, “the Trail of Tears”, “D-Day”, “The Declaration of Independence Signing”, “the Nakba”, “the Napoleonic Wars”
  • $\bullet$
    Politics
    • $\circ$
      Ideological orientations: “a Republican”, “a Democrat”, “a rightist”, “a leftist”, “a centrist”, “a Trump supporter”, “a Harris supporter”, “a libertarian”, “a progressive”, “a nationalist”, “an anarchist”, “a socialist”, “a capitalist”, “a communist”, “a fascist”, “an authoritarian”
    • $\circ$
      Current conflicts and issues:
        Geopolitical: “the Israel-Palestine conflict”, “the Ukraine-Russia war”, “a protester in Hong Kong”, “a Kurdish fighter”, “a Syrian refugee”
        Gun Rights and Policing: “a mass shooter”, “a gun owner”, “a gun control advocate”, “a mass shooting survivor”, “a police officer in a high-crime area”, “a police brutality victim”
        Immigration and Borders: “a migrant crossing the US border”, “a border patrol agent”
A.3. Full Survey Questions
We ask users to provide responses on a 5-point Likert-style scale (1 $=$ “not at all”, 3 $=$ “somewhat”, 5 = “entirely”) after interacting with each interface for each of the following statements:
  • Satisfaction: “I am satisfied with the collage I created”
  • Rethinking: “Interacting with the interface made me rethink the mental picture of the subject I had right before using this interface”
  • Appropriateness: “I thought the content generated from the interface (e.g. images, suggestions) was appropriate to the prompt”
  • Control: “I feel like the interface gives me detailed control over the image generation process”
  • Interest: “I found the content of the suggestions interesting”
At the end of the study, participants were also asked to rank each of the interfaces by agreement with the rethinking, appropriateness, and control statements. This provided another data signal as well as a moment for participants to reflect upon the entire interview and make open-ended comments. Participants were allowed to give two interfaces the same rank if there were strong ties.
Appendix B Detailed View of Interfaces
B.1. Design Iteration for Agonistic
In the first iteration of Agonistic, we generated different set of suggestions for each subject in the user’s prompt (an example with only one subject is shown in 7). Finding that generating suggestions for multiple subjects was too costly and slow for interview purposes, we stopped organizing suggestions by subject and instead only generated suggestions for the main subject of the prompt, as shown in 1.
We also asked users to input their “mental image” of their subject before generating images in the initial iteration of Agonistic, for the purpose of generating suggestions that challenged the user’s mental image. Since we found this feature to be prohibitively difficult to use when regenerating suggestions or changing prompts, however, we removed it in favor of using LLM-generated mental image descriptions for the same purpose, as detailed in §3.
Figure 7. First iteration of Agonistic interface
B.2. Diverse
We use GPT-3.5 with the full alleged leaked DALL-E system prompt (Milmo and Kern, 2024) listed below to rewrite the user’s prompt four times before passing each rewritten prompt into the image generation model.
You are given the prompt ”prompt.” Diversify depictions of people in this prompt to always include always DESCENT and GENDER for EACH person using direct terms. Adjust only human descriptions. EXPLICITLY specify these attributes, not abstractly reference them. The attributes should be specified in a minimal way and should directly describe their physical form. Your choices should be grounded in reality. For example, all of a given OCCUPATION should not be the same gender or race. Additionally, focus on creating diverse, inclusive, and exploratory scenes via the properties you choose during rewrites. Make choices that may be insightful or unique sometimes. Use ”various” or ”diverse” ONLY IF the description refers to groups of more than 3 people. Do not change the number of people requested in the original description. Don’t alter memes, fictional character origins, or unseen people. Maintain the original prompt’s intent and prioritize quality. Do not create any imagery that would be offensive. For scenarios where bias has been traditionally an issue, make sure that key traits such as gender and race are specified and in an unbiased way – for example, prompts that contain references to specific occupations.
B.3. Reformulative
Users begin by entering their prompt in the prompt box at the top of the interface.
When they press enter or click on the $\gg$ symbol on the prompt box, reformulations are generated and displayed to the user in the “Suggestions” box to the left of the interface. Each reformulation includes a thumbnail to visualize how the reformulated prompt might be visualized.
When users click on a suggestion, a new “Prompt Workspace” pane opens. In this pane, users can edit the reformulation in a text box. Clicking the “Generate images” button generates images using the prompt in the text box. The images are persistent and stay on the panel even when the user clicks the $\gg$ button in the Prompt Workpace Pane and returns to the Suggestions list.
B.4. Agonistic
Users begin by entering their prompt in the prompt box at the top of the interface.
When they press enter or click on the $\gg$ symbol on the prompt box, interpretations are generated and displayed to the user in the “Possible Interpretations” box to the left of the interface. Each card includes a decription of the interpretation, a source including the Wikipedia page title and section, and a thumbnail.
When users click on a card, it expands and a justification of the form “You may assume [X], but [Y]” is displayed, along with a clickable link to the Wikipedia page and section the interpretation references. After three seconds, an “Accept” button appears.
Clicking the “Accept” button opens the Prompt Workspace pane. In this pane, users can edit the reformulation in a text box. Clicking the “Generate images” button generates images using the prompt in the text box. The images are persistent and stay on the panel even when the user clicks the $\gg$ button in the Prompt Workpace Pane and returns to the Possible Interpretations list.
Appendix C Inter-Rater Reliability Methodology
We calculated inter-rater reliability (IRR) on value codes. the two first co-authors first independently coded a subset of 3 participant interviews using the coding ontology described in §4.3. Since some codes were not mutually exclusive with each other, we calculated a separate Cohen’s kappa score for each code, with the following results: 0.64 (Direct), 0.00 (Reminder), 0.65 (Expansion), 1.00 (Challenge), 0.66 (Realism), 0.83 (Familiarity), 0.60 (Diversity), 0.38 (Aesthetics) We then took an average of Cohen’s kappa scores across value codes, weighted by the frequency of each value code, to calculate our final IRR of 0.67.
We noticed that the Cohen’s kappa score for several codes might be unreliable to the low frequency of certain codes in the subset of interviews used to calculate IRR (for instance, there were only 3 instances of Reminder  and 1 instance of Challenge ). This issue was especially severe for intent codes, due to the extremely high proportion of Direct codes compared to other intents. For this reason, two first co-authors then reviewed all non-Directintent codes together to establish greater agreement. We note that while this approach reduced the chance of false positives for non-Directintent codes, it did not reduce the chance of false negatives. Therefore, the amount of reflection captured by our coding in this study is a conservative estimate of the amount of reflection that may have occurred.
Appendix D Extended Results
D.1. Rankings for change in mental image
At the end of the study, participants are asked to rank the interfaces by how much each made them rethink their mental picture (compared to right before using it), allowing equal ranks between strongly tied interfaces. This ranking allows participants to rank the interfaces with the benefit of retrospection (e.g., participants may not realize how interacting with an interface changed their mental picture until later, the scale may have been disrupted) and provides another view into reflection. Agonistic was ranked higher than Reformulative by $79\%$ of participants, Diverse by $86\%$, and Baseline by $89\%$. Reformulative was ranked higher than Diverse by $70\%$ and Baseline by $68\%$; Diverse higher than Baseline by $52\%$. This suggests that participants often believe Reformulative made them rethink their mental picture, although they may not have realized this at the moment. Diverse and Baseline are ranked above each other by $50\%$ of participants and therefore are overall similar in retroactive attribution towards reflection.
D.2. Interface Ordering Analysis
[…]
D.3. Change in design statement
Participants write an “design statement” — one or two sentences describing the most important aspects of the subject represented in the collage – after producing an initial collage with Baseline. The design statements provided both a more standardized and a richer way to record the progression of participant assumptions over different interfaces beyond the data collected by the self-reported rankings. After iterating through each following interface, participants copied over their previous design statement and made modifications to reflect the changes to their collage. They could opt to make no changes even if images were replaced, should they feel that the previous design statement satisfactorily described the new collage. The design statement reflects a different dimension of reflection than either the change in mental image or how images are added: for example, some participants did not report a substantive changes in mental picture of the subject but still modified the design statement to reflect a change in how they viewed their collage, like by emphasizing some aspects of the collage; some participants also replaced images in the collage but did not change the design statement because they felt that it applied just as well (the changes fell within the scope of the previous design statement). The design statement uniquely captures the self-perception of the participant as a collage designer synthesizing their choices.
Interface Levenshtein Embeddings
Raw Scaled Raw Scaled
Diverse  6.72 0.27 0.02 0.28
Reformulative 9.24 0.38 0.05 0.32
Agonistic 12.31 0.57 0.08 0.56
Table 3. Mean change in participant design statement after using each interface, as measured by Levenshtein and embedding-based distance. “Scaled” indicates 0-1 min-max scaling, where the smallest change is mapped to 0 and the highest is mapped to 1.
We can measure reflection under each interface via change in design statement by computing the mean text distance after using an interface. One purely syntactic metric is Levenshtein distance across words, which measures the number of insertions, deletions, and additions needed to transform one text to another. Another metric that captures more semantic information is to compare the cosine similarity between the BERT representations of the text samples. In each instance, Agonistic has a higher level of change than Diverse ($p\ll 0.05$ except for Levenshtein raw, where $p=0.06$). The relationship between Reformulative and either Diverse or Agonistic is not statistically significant, except for scaled embeddings, in which Agonistic is larger than Reformulative with $p=0.04$.
D.4. Example Collage Progression
Figure 8 shows an example of a collage progression from P6 for the subject “Jesus”. Observe that under Diverse, the participant adds images featuring demographically diverse individuals worshipping Jesus. After previously choosing not to include a darker-skinned picture of the subject “Jesus” produced by Diverse, the participant was presented with two interpretations of Jesus by Agonistic— Jesus as an olive-skinned Jew and as a red-haired man (from Islamic records) — both of which they ended up accepting after much thought and reading the respective Wikipedia pages. They reflected: “It’s really interesting knowing there are so many interpretations of who Jesus is, I think I only thought about how I got taught in the Bible at my Church, but I never really thought about other people think about who Jesus is.” Under Reformulative, the participant added several images that were closer to what they originally imagined Jesus to look like, notably with two of the images in an illustration style.
(a) Baseline
(b) Diverse
(c) Agonistic
(d) Reformulative
Figure 8. Example collage progression (left-to-right). Images added/replaced in each new round are fully shown; kept images are faded.
D.5. Table Form for Results
Interface 5-Point Rating Min-Max Scaled
Overall Identity Politics History
Baseline $2.03\pm 1.00$ $1.90\pm 1.14$ $2.20\pm 0.87$ $2.00\pm 0.94$ $0.43\pm 0.44$
Diverse  $1.48\pm 0.93$ $1.50\pm 1.02$ $1.60\pm 1.02$ $1.33\pm 0.67$ $0.17\pm 0.35$
Reformulative $1.93\pm 1.05$ $1.80\pm 0.98$ $2.60\pm 1.11$ $1.33\pm 0.47$ $0.37\pm 0.41$
Agonistic $2.97\pm 1.19$ $2.90\pm 0.94$ $3.00\pm 1.18$ $3.00\pm 1.41$ $0.78\pm 0.39$
Table 4. Comparison of overall, identity, politics, history prompt ratings, and min-max scaled scores across interfaces (mean $\pm$ standard deviation). Min-max scaling is applied to each participant, with their lowest rating set to 0 and their highest set to 1. Visualized in Figure 2.
Interface Rethink Appropriateness Control Interestingness
Baseline $2.03\pm 1.00$ $3.31\pm 0.91$ $3.03\pm 1.22$
Diverse  $1.48\pm 0.93$ $3.07\pm 1.20$ $2.59\pm 1.16$
Reformulative $1.93\pm 1.05$ $3.93\pm 0.91$ $3.79\pm 1.21$ $4.00\pm 1.11$
Agonistic $2.97\pm 1.19$ $3.83\pm 1.15$ $3.90\pm 0.99$ $4.28\pm 0.91$
Table 5. Comparison of rethink, appropriateness, control, and interestingness (mean $\pm$ standard deviation) across interfaces. Visualized in Figure 4.
Interface Values
Realism Familiarity Diversity Aesthetics
Baseline $0.40$ $0.44$ $0.26$ $0.08$
Diverse  $0.42$ $0.47$ $0.24$ $0.05$
Reformulative $0.24$ $0.53$ $0.38$ $0.10$
Agonistic $0.39$ $0.42$ $0.45$ $0.08$
Table 6. Comparison of the proportion of images added under each value, across interfaces. Rows do not sum to 1 because values are non-exclusively coded. Visualized in Figure 6.
Interface Intents
Direct Reminder Expansion Challenge
Baseline 0.98 $0.00$ $0.02$ $0.00$
Diverse  0.95 $0.00$ $0.05$ $0.00$
Reformulative 0.90 $0.04$ $0.06$ $0.00$
Agonistic 0.67 $0.11$ $0.19$ $0.04$
Table 7. Comparison of the proportion of images added with each type of intent, across interfaces. Rows sum to 1. Visualized in Figure 4.
Values Intents
Direct Reminder Expansion Challenge
Realism 0.879 $0.047$ $0.065$ $0.009$
Familiarity 0.954 $0.031$ $0.015$ $0.000$
Diversity 0.753 $0.067$ $0.157$ $0.022$
Aesthetics 0.960 $0.000$ $0.040$ $0.000$
Table 8. Comparison of the proportion of ratings across values and intents. Rows sum to 1. Visualized in Figure 6.
D.6. Quantitative Coding Analysis
A notable difference between the quantitative analysis of intent code occurrences and 5.1 is the finding that Diverse induces greater degrees of ‘Reminder’ and ‘Expansion’ of intent than Baseline ($+0.02$ and $+0.04$, respectively), in contrast to self-reported findings that Baseline induces greater reflection than Diverse. These differences may be explained by the fact that coding for added images excludes some forms of reflection, such as reflection that does not explicitly correlate to a particular image or reflection caused by images that participants ultimately reject, as further discussed in §6. For this reason, the methodology of this section may more accurately capture the unique reflection due to each interface because it gives less weight to reflection from external causes like performing the task itself. The ‘hierarchy’ of interfaces suggested by our findings in this section is thus $\left[\textsf{{Agonistic}}~{}\gg\textsf{{Reformulative}}~{}>\{\textsf{{Baseline}},\textsf{{Diverse}}\}\right]$.
D.7. Additional Qualitative Themes
Agonistic can feel like a double-edged sword, treading the line between appropriate and inappropriate. The results in §6.1 suggest that appropriateness is correlated with higher rethinking. However, by design, Agonistic often presents users with interpretations that are perceived initially by users as unrelated, inappropriate, or even offensive. For many participants who underwent reflection, these interpretations eventually seemed appropriate, explaining the high appropriateness rating of Agonistic. However, for other participants, these interpretations remained inappropriate (even after reflection). Several participants commented on the dually appropriate and inappropriate character of Agonistic. P4 suggested that “it [Agonistic] was giving me the most appropriate, but it was also giving me the most inappropriate [outputs] at the same time… [it] would sometimes hit a home run, sometimes wildly off track.’P16 said that Agonistic was “good good and bad bad […] I like the specificity [of Agonistic]. I just think the specificity was kind of whacked out in a way where it just wasn’t super on the bar of what I was thinking about”. Together, these suggest that the experience of using Agonistic can vary widely, and that perhaps designing for reflection via agonism means that interface outputs will not seem appropriate to many users by its own nature.
Appendix E Extended Discussion
E.1. The Many Heads of Reflection
We find it important to note that people find reflection across all the interfaces, as demonstrated by Table 1; although the overall occurrence of reflection does vary substantively (as discussed in §5), reflection is still a very subjective, personal experience. Indeed, while Baseline was uninspiring for many — “I typed what I thought and I got what I thought I would getP21 – it was a highly reflective experience for others — “I have a lot of other potential images in my mind of it now, and I don’t know what is […] actually trueP28. We find that reflection generally seems to emerge from an initially unexpected but eventually rationalizable encounter with the interface (either the images themselves or contextualizing content). This reflects a host of work on reflection and related cognitive activities emphasizing encounter with alterity or the Other (Ye et al., 2024; Halbert and Nathan, 2015; Levinas, 1969; Lacan, 2006, inter alia). Although reflection often occurred in planned ways (i.e., as a result of design elements we had introduced with the intention of inducing reflection), it also often emerged in spontaneous but nonetheless personally significant ways. This emphasizes that image generation interfaces may benefit from more adaptive design, but more importantly that users will always reflect in ways that interface designers cannot completely control.